OpenStudy (anonymous):

Nina makes and sells necklaces. For n necklaces, her revenue function is r(n) = 11.75n and her cost function is c(n) = 8.9 + 1.2n. Which function is Nina’s profit function? p(n) = 10.55n – 8.9 p(n) = 12.95n – 8.9 p(n) = 12.95n + 8.9 p(n) = 14.1n + 8.9

OpenStudy (dumbcow):

profit = revenue - cost

OpenStudy (anonymous):

@dumbcow I'm still confused. @sammixboo @wio can you guys help please?

OpenStudy (jingle535):

Crazy to figure out which one is the profit function you need to subtract the cost function c(n) = 8.9 + 1.2n by the revenue function r(n) = 11.75n . So r(n) = 11.75n - c(n) = 8.9 + 1.2n is...

OpenStudy (anonymous):

\[ p(n) = r(n)-c(n) = (11.75n )- (8.9 + 1.2n) \]

OpenStudy (jademc01):

p(n) = 10.55n – 8.9